Calculate the volume, in mL, of 1.72 M HCl solution that will react with 2.67 mol of CaCO3?


1
Expert's answer
2022-03-24T05:24:02-0400

CaCO3 + 2HCl ----- CaCl2 + H2CO3

From the equation,

1 mol of CaCO3 reacts with 2 mol of HCl

2.67 mol of CaCO3 will react with (2.67*2) mol of HCl = 5.34 mol of HCl


Molarity = mol/volume in L

1.72 = 5.34/V

V = 5.34/1.72 = 3.10L = 3100mL